OpenStudy (anonymous):

Find all radian measures of A. (2sinAcosA)^2+4sin2A-1=0

OpenStudy (mathmath333):

use \(\large \color{black}{\begin{align}2\sin a \cdot \cos a=\sin 2a\hspace{.33em}\\~\\ \sin2a=x\hspace{.33em}\\~\\ \end{align}}\) then solve the quadratic

OpenStudy (anonymous):

Do I plug in 2sina*cosa in to sin2a?

OpenStudy (mathmath333):

substitute first \(\large \color{black}{\begin{align} \sin2a\hspace{.33em}\\~\\ \end{align}}\) for \(\large \color{black}{\begin{align} 2\sin a\cdot \cos a\hspace{.33em}\\~\\ \end{align}}\)

OpenStudy (anonymous):

i did that. I got (sin2A)^2+4sin^2A-1=0

OpenStudy (mathmath333):

ok now substitute \(x\) for \(\sin 2a\)

OpenStudy (anonymous):

ok I did, then what do I do?

OpenStudy (mathmath333):

you will get this equation \(\large \color{black}{\begin{align} x^2+4x-1=0\hspace{.33em}\\~\\ \end{align}}\) u can now apply the quadratic formula to find \(x\) or use completing the square method

OpenStudy (anonymous):

so after using the quadratic formula i got 0.23 radians and 2.9 radians

OpenStudy (anonymous):

but the correct answer is 0.12 and 3.02

OpenStudy (mathmath333):

after solving the quadratic you will get you will get this equation \(\large \color{black}{\begin{align} x=0.23\hspace{.33em}\\~\\ x=-4.23\hspace{.33em}\\~\\ \end{align}}\) now substitute back \(\large \color{black}{\begin{align} \sin 2a=x\hspace{.33em}\\~\\ \end{align}}\)

OpenStudy (anonymous):

ok thank you